OpenStudy (anonymous):

Assuming 100% dissociation, calculate the freezing point and boiling point of 2.05 m K3PO4(aq). Constants may be found here.

OpenStudy (aaronq):

use the formula: \(\Delta T=i*m*K_{\theta}\) where, \(\theta=b\) when finding BP, and \(\theta=f\) when finding freezing point (By, that i mean use the appropriate constant.) More on this: http://openstudy.com/study#/updates/521ecb3ae4b0750826e0c362